Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 286     Entry time: 04 Aug 2023, 07:10
Author: Giuseppe Mercurio, Teguh Asmara 
Group: SCS 
Subject: night shift summary 

achievements:

  • changed the LEPM to optimize resolution
  • found 140 meV

problems:

  • were not able to recover settings found because it is difficult to reproduce LEPM positions

plan:

  • try to improve resolution
ELOG V3.1.4-7c3fd00